OpenStudy (jessicawade):

does this converge or diverge? If the sequence does converge, what number do the terms of the sequence converge to?

OpenStudy (jessicawade):

\[a^n=(-1)^n100*n \over 100+n\]

TheSmartOne (thesmartone):

\(\Large a^n=\)\(\Large \frac{(-1)^n100*n}{100+n}\)

OpenStudy (jessicawade):

yes

OpenStudy (jessicawade):

do you knwo?

TheSmartOne (thesmartone):

I would say diverges... not sure though.. :/

OpenStudy (jessicawade):

oh hmm

TheSmartOne (thesmartone):

the last example is pretty helpful... http://www.wolframalpha.com/widgets/view.jsp?id=a787670f0f1047d7fbe288763c55ba14

TheSmartOne (thesmartone):

so i would say Divierge @jessicawade

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne the last example is pretty helpful... http://www.wolframalpha.com/widgets/view.jsp?id=a787670f0f1047d7fbe288763c55ba14 \(\color{blue}{\text{End of Quote}}\) Referring to the last example in Khan's video.

OpenStudy (perl):

it does not converge because of the (-1)^n , it alternates. The absolute terms of the sequence converge to 100, but the (-1)^n makes it alternate so for very large n you have the sequence ... , -99.99, 99.999, -99.9999, etc

TheSmartOne (thesmartone):

Yup exactly like the last example in Khan's video I gave :P